Graph the following linear equation by finding x- and y-intercepts.
2x – 4y = -8

Describe the graph verbally by showing the calculation of the two intercepts and giving the ordered pair for each. Describe the line you have graphed. The description should include the quadrants through which the line passes, and the direction the line leans (forward or backward).


Found 3 solutions by Tobiasz, Alan3354, richwmiller:
Answer by Tobiasz(54)

Nice Question:
In order to find the x or y value you most put zero(0) as a input as that x or y variable.
Lets begin to find x__+first
2x - 4y = -8
2x - 4(0) = -8
2x = -8
divide 2 from both sides
x = -4 <--- one coordinate
lets find y now.
2(0) -4y = -8
-4y = -8
divide -4 from both sides and you get:
y =2
(-4,2)






Answer by Alan3354(69443)

Graph the following linear equation by finding x- and y-intercepts.
2x – 4y = -8
-------------------
set x = 0 to find the y-intercepts, and vice versa
x - 2y = -4 (I reduced it)
x = 0: -2y = -4
y = 2
--> (0,2) y-intercept
------------------
y=0: x = -4
--> (-4,0) x-intercept
Draw a line thru the 2 points

Answer by richwmiller(17219)

2x-4y=8
let x=0 and find y
(0,y)
0-4y=8
y=-2
(0,-2)
let y= 0 and find x
2x=8
x=4
(4,0)
so our two ordered pairs are (0,-2) and (4,0)
To find which way the line leans we need to find the slope
we can do it two ways
y-y/x-x=m
or put the equation in the form y=mx+b where m is the slope
since we went through the bother of finding the two points lets use them
y-y/x-x=m the slope
-2-0/0-4=-2/-4=1/2
so the line slants or leans up from left to right
so when x=0 y is negative but rises to zero when x=4